1) 1.00 L flask is filled with 1.30 g of argon at 25 ∘C. A sample of ethane vapor is added to the same flask until the total pressure is 1.200 atm

A) What is the partial pressure of argon, PAr, in the flask?

b)What is the partial pressure of ethane, Pethane, in the flask?
Answer #1

mol of argon = masS/MW = 1.3/39.9 = 0.03258 mol of Ar

partial pressure of argon

P V= nRT

P = nRT/V = 0.03258*0.082*298/1

P = 0.7961 Atm (partial pressure of Argon)

Pt = Peth + Par

1.2 =Peth + 0.7961

Peth = 1.2-0.7961 = 0.4039 atm